THE ν_{1} and ν_{2} BANDS OF DOBr AND THE ROTATIONAL SPECTRA OF THE v_{2} = 1 AND v_{3} = 1 STATES

\underline{E. A. Cohen}, H. S. P. Müller, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California 91109; T. L. Tan, Department of Physics, Faculty of Science, National University of Singapore, Lower Kent Ridge Road, Singapore 0511, Singapore; G. A. McRae, Atomic Energy of Canada Limited, Chalk River Laboratories, Chalk River, Ontario K0J 1J0, Canada; J. W. C. Johns,, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada; and M. Noël, Institute for National Measurement Standards, National Research Council of Canada, Ottawa, Ontario K1A 0R6, Canada.

The ν_{1} and ν_{2} bands of DOBr centered near 2673.2 and 853.4 cm^{−1} respectively have been observed at 0.006 cm^{−1} resolution. The ν_{1} band is perturbed by an anharmonic resonance (ΔK_{a} = 0) with 3ν_{3}+ν_{2}. In addition, the millimeter spectra arising from the v_{2} = 1 and v_{3} = 1 states have been observed. All rotational and vibrational spectra from both bromine isotopic species have been fitted with a single calculation. The perturbation in the ν_{1} band has been well described. Equilibrium rotational and centrifugal distortion constants have been determined. The equilibrium structure has been derived from the DOBr and HOBr rotational constants. The harmonic force field has been calculated and will be compared with that of HOCl and with those derived from ab initio calculations. Changes in quadrupole coupling with the BrO stretch and DOBr bend have been determined.